OpenStudy (anonymous):

Possible topics for discussion include: •Absolute Value •Inequalities •Literal and Linear Equations •Variation •Functions •Polynomial Division •Factoring

OpenStudy (anonymous):

can tell what they are please

OpenStudy (anonymous):

isn't this math hehe

OpenStudy (anonymous):

absolute values is the exact value of things like the absolute value of |-3 | is 3

OpenStudy (anonymous):

inequalities are the equalities that lack to be equal like 3>2 or 4<5

OpenStudy (anonymous):

linear equations hmmm well this its a bit rough for me to remember but I believe its a few points that when you graph the points it makes a perfect line in the graph (3,4)(5,4)(1,4) or (4,1)(4,3)(4,5) as long as you plot the points on a graph and it makes a stait line your good to go

OpenStudy (anonymous):

Variation okay I don't remember this one but heres a page that cam help http://www.mathsisfun.com/data/standard-deviation.html

OpenStudy (anonymous):

A function is a special relationship between values: Each of its input values gives back exactly one output value. It is often written as "f(x)" where x is the value you give it. Example: f(x) = x/2 ("f of x is x divided by 2") is a function, because for every value of "x" you get another value "x/2", so: * f(2) = 1 * f(16) = 8 * f(-10) = -5